# Non-commutative effective field theory of the lowest Landau level superfluid

### v2

@@ V Discussion
We have derived the non-commutative EFT of the LLL superfluid starting from a Hamiltonian for bosons restricted to the LLL, using a coherent state path integral approach. Although this work does not produce any qualitatively new results about this phase, it gives insight into the origin of the non-commutative fields which was lacking in previous works. Unlike the previous work [1] which justifies this non-commutative EFT using boson-vortex duality, this direct approach has the advantage of being able to make quantitative statements about the low-energy physics and comparison to the microscopic parameters of the model.
@@
A related aspect we want to emphasize is that this derivation suggests that the EFT for the LLL superfluid discussed here and in Ref. 1 is not an NC field theory in the sense that it is used in older literature [18]. One of the hallmarks of NC field theory is the presence of UV-IR mixing [52], where an IR length scale $L$ is related to a UV length scale $l_{B}^{2}/L$. This feature often manifests in lowest Landau level theories such as HLR theory as a relationship between the momentum of a quasiparticle and its dipole moment [22]. In contrast, the theory developed in this paper has a UV distance cutoff $\Lambda^{-1}$ that is greater than the length scale set by the non-commutativity parameter, and therefore it cannot capture the effects of such UV-IR mixing. Nevertheless, it is important to appreciate that in this context the NC field theory arises as a natural language to keep track of the symmetries for the EFT, and is unlike UV complete non-commutative field theory like the one suggested for the HLR state and its descendants [19, 21]. On the formal side, it may be appropriate to think of two qualitatively different kinds of NC field theories, parametrized by the dimensionless ratio between the two characteristic length scales of the EFT – the UV cutoff $\Lambda^{-1}$ and the non-commutativity parameter $l_{B}$.
@@
Exploring the dynamical consequences of this NC field-theoretic description is an interesting direction for future work, particularly hydrodynamics in the presence of slowly varying trap potentials. A varying potential breaks magnetic translation symmetry and thus dipole conservation [42], leading to chiral charge transport perpendicular to the potential gradient. A similar mode should also emerge at the boundary of rotating BECs [53]. A quantitative analysis of this mode within the NC effective field theory framework is left for future work.
@@
Another open problem is the physics of vortex lines of $\hat{\varphi}$ in space-time [44, 45]. Large vortex loops of the coarse-grained field $\hat{\varphi}$ confined to a time slice correspond to cooperative ring exchange processes [54, 55, 56] in the triangular vortex lattice (since the system lacks Lorentz invariance, space and time are not on equal footing). Reference 57 suggests, based on an approximate calculation using vortex coordinates, that the proliferation of cooperative ring exchange processes in the path integral would drive a transition to a bosonic fractional quantum Hall state. It would be valuable to reproduce their results using the effective field theory developed in this work.
@@
We also note a related study on LLL superfluid hydrodynamics [58], which employs a different coherent-state path integral approach to describe rotating BECs. Unlike our work, which focuses on vortex lattices and their fluctuations, their approach describes experimentally relevant cases where the BEC condenses into a single angular momentum state.
@@ Appendix B Tools in non-commutative field theory
We describe the useful mathematical tools in non-commutative field theory in this appendix. Most of this discussion can be found in standard reviews of non-commutative field theory, such as [18].
@@ C.1 Estimating $g^{\ast}$
Note that a more efficient computation of the factor $1.160$ is described in [40].

### v3

@@ V Discussion
We have derived the non-commutative EFT of the LLL superfluid starting from a Hamiltonian for bosons restricted to the LLL, using a coherent state path integral approach. Although this work does not produce any qualitatively new results about this phase, it gives insight into the origin of the non-commutative fields which was lacking in previous works. Unlike the previous work [Du2024a] which justifies this non-commutative EFT using boson-vortex duality, this direct approach has the advantage of being able to make quantitative statements about the low-energy physics and comparison to the microscopic parameters of the model.
@@
A related aspect we want to emphasize is that this derivation suggests that the EFT for the LLL superfluid discussed here and in Ref. Du2024a is not an NC field theory in the sense that it is used in older literature [DouglasNekrasov2001]. One of the hallmarks of NC field theory is the presence of UV-IR mixing [Minwalla2000], where an IR length scale $L$ is related to a UV length scale $l_{B}^{2}/L$. This feature often manifests in lowest Landau level theories such as HLR theory as a relationship between the momentum of a quasiparticle and its dipole moment [PasquierHaldane1998]. In contrast, the theory developed in this paper has a UV distance cutoff $\Lambda^{-1}$ that is greater than the length scale set by the non-commutativity parameter, and therefore it cannot capture the effects of such UV-IR mixing. Nevertheless, it is important to appreciate that in this context the NC field theory arises as a natural language to keep track of the symmetries for the EFT, and is unlike UV complete non-commutative field theory like the one suggested for the HLR state and its descendants [DongSenthil2020, Goldman2022]. On the formal side, it may be appropriate to think of two qualitatively different kinds of NC field theories, parametrized by the dimensionless ratio between the two characteristic length scales of the EFT – the UV cutoff $\Lambda^{-1}$ and the non-commutativity parameter $l_{B}$.
@@
Exploring the dynamical consequences of this NC field-theoretic description is an interesting direction for future work, particularly hydrodynamics in the presence of slowly varying trap potentials. A varying potential breaks magnetic translation symmetry and thus dipole conservation [Spodyneiko2023], leading to chiral charge transport perpendicular to the potential gradient. A similar mode should also emerge at the boundary of rotating BECs [Jeevanesan2022]. A quantitative analysis of this mode within the NC effective field theory framework is left for future work.
@@
Another open problem is the physics of vortex lines of $\hat{\varphi}$ in space-time [NguyenMoroz2023, Du2024b]. Large vortex loops of the coarse-grained field $\hat{\varphi}$ confined to a time slice correspond to cooperative ring exchange processes [KivelsonCRE1, KivelsonCRE2, LeeBaskaranKivelsonCRE] in the triangular vortex lattice (since the system lacks Lorentz invariance, space and time are not on equal footing). Reference GhoshBaskaranCRE suggests, based on an approximate calculation using vortex coordinates, that the proliferation of cooperative ring exchange processes in the path integral would drive a transition to a bosonic fractional quantum Hall state. It would be valuable to reproduce their results using the effective field theory developed in this work.
@@
We also note a related study on LLL superfluid hydrodynamics [MusserGoldmanSenthil2024], which employs a different coherent-state path integral approach to describe rotating BECs. Unlike our work, which focuses on vortex lattices and their fluctuations, their approach describes experimentally relevant cases where the BEC condenses into a single angular momentum state.
@@ Appendix B Tools in non-commutative field theory
We describe the useful mathematical tools in non-commutative field theory in this appendix. Most of this discussion can be found in standard reviews of non-commutative field theory, such as [DouglasNekrasov2001].
@@ C.1 Estimating $g^{\ast}$
Note that a more efficient computation of the factor $1.160$ is described in [YangMacdonald2004].